# Seeing What’s Not There: Spurious Correlation in Multimodal LLMs

### v1

@@ Abstract
Unimodal vision models are known to rely on spurious correlations, but it remains unclear to what extent Multimodal Large Language Models (MLLMs) exhibit similar biases despite language supervision. In this paper, we investigate spurious bias in MLLMs and introduce SpurLens, a pipeline that leverages GPT-4 and open-set object detectors to automatically identify spurious visual cues without human supervision. Our findings reveal that spurious correlations cause two major failure modes in MLLMs: (1) over-reliance on spurious cues for object recognition, where removing these cues reduces accuracy, and (2) object hallucination, where spurious cues amplify the hallucination by over 10x. We validate our findings in various MLLMs and datasets. Beyond diagnosing these failures, we explore potential mitigation strategies, such as prompt ensembling and reasoning-based prompting, and conduct ablation studies to examine the root causes of spurious bias in MLLMs. By exposing the persistence of spurious correlations, our study calls for more rigorous evaluation methods and mitigation strategies to enhance the reliability of MLLMs.
@@ 1 Introduction
Object: Tie, Spurious Cue: Suit
@@
Object: Toaster, Spurious Cue: Kitchen Counter
@@
There is a black coffee maker on the countertop, but I don’t see a toaster in the image. The area appears to have kitchen…
Object: Balance Beam, Spurious Cue: Rope
@@
Object: Ruffed Grouse, Spurious Cue: Moss
@@
Figure 1: Some failures of GPT-4o-mini (accessed in March 2025) identified by our method. (Left) The model fails to recognize objects in the absence of spurious cues. (Right) Spurious cues trigger hallucinations.
Multimodal large language models (MLLMs) [32, 12, 16, 22] have seen rapid advances in recent years. These models leverage the powerful capabilities of large language models (LLMs) [23, 30] to process diverse modalities, such as images and text. They have demonstrated significant proficiency in tasks such as image perception, visual question answering, and instruction following.
However, despite these advancements, MLLMs still exhibit critical visual shortcomings [28, 29]. One such failure is object hallucination [8, 2, 14, 7], where MLLMs generate semantically coherent but factually incorrect content, falsely detecting objects that are not present in the input images, as exemplified in Fig. 1. We hypothesize that many of these failures stem from a well-known robustness issue in deep learning models: spurious bias – the tendency to rely on non-essential input attributes rather than truly recognizing the target object [35]. While spurious correlation reliance has been extensively documented in single-modality image classifiers, the extent to which it persists in MLLMs remains unclear.
Understanding spurious correlations is crucial because they can lead to systematic failures. Consider the case of a fire hydrant (Fig. 2). When a fire hydrant appears in a street scene, Llama-3.2 [16] correctly recognizes it 96% of the time. However, when placed in unusual contexts, such as a warehouse, accuracy drops to 83%, suggesting an over-reliance on contextual cues rather than the object itself. Conversely, when shown a street scene without a fire hydrant, the model hallucinates its presence 7 times more often than when given a random image without a hydrant, such as an indoor environment lacking street-like features.
Our results reveal that MLLMs over-rely on spurious correlations, leading to two major failures. The first is object recognition failures (false negatives) – when spurious cues are removed, recognition performance drops significantly. For example, GPT-4o-mini [22] sees a 17.9% accuracy drop on COCO [9], from 87.6% to 69.7% (Tab. 1). We illustrate some of these failures in Fig. 1. The second is object hallucination (false positives) – spurious cues trigger hallucinations, causing models to misidentify objects. For GPT-4o-mini, hallucination rates increase 15-fold when spurious cues are present (Tab. 2). We provide some qualitative examples of these hallucinations in Fig. 1.
@@
SpurLens offers three key advantages over prior work. First, SpurLens enables the detection of a broader range of spurious cues, as it is not limited to manually predefined features. Second, it provides natural language descriptors for detected cues, offering insight into model biases. Finally, our method requires no human supervision, making it scalable and adaptable across datasets and object classes.
@@
While SpurLens allows us to systematically analyze spurious correlations in MLLMs, some key questions remain: from where do these biases originate, and can they be mitigated? To address these, we conduct a series of ablation studies, each designed to isolate different factors contributing to spurious bias. First, we investigate whether spurious bias persists even when all visual evidence of an object is artificially removed. We do this by dropping the corresponding visual tokens before they reach the model’s vision encoder. We find that even when target objects have been excised from images, MLLMs tend to hallucinate their presence at a high rate. Next, we explore whether alternative prompting strategies can mitigate spurious reliance. We find that prompt ensembling, where we take th majority vote between queries with different question phrasing, provides slight improvements. We also examine reasoning-based prompting, where the model first describes the image before answering, but observe a trade-off: while it reduces hallucination rates, it also lowers object recognition accuracy, indicating that it does not fully resolve spurious bias. Finally, we examine whether spurious correlations can be measured purely from the MLLM image embeddings. This helps determine whether these biases originate from visual processing or from the interaction between vision and language components. We find that the vision encoder alone also exhibits spurious biases, reinforcing that the issue extends beyond multimodal fusion errors. These findings indicate that spurious bias is a deeply rooted problem in MLLMs and that simple mitigation strategies are insufficient to fully address it.
Is there a fire hydrant in the image?
With Spurious Cue
Without Spurious Cue
Object Present
Refer to caption Yes
Refer to caption No
Object Absent
Refer to caption Yes
Refer to caption No
Figure 2: (Top) Object Recognition - the model overrelies on spurious cues for recognition. (Bottom) Object Hallucination - spurious cues amplify hallucinations.
@@ 2 Related Work
Spurious Correlation: Spurious correlations have been extensively studied in the context of deep neural network classifiers (e.g., ViT [1]), with various approaches proposed to detect and mitigate the issue [26, 5, 21]. However, these studies primarily focus on uni-modal settings (image classification tasks). Some research [33, 31, 4] has explored spurious correlations in CLIP [24], framing the problem in terms of zero-shot performance across vision and language modalities. Ye et al. [35] introduces a visual question answering (VQA) benchmark designed to evaluate MLLMs’ reliance on spurious correlations using open-source image datasets. Zheng et al. [37] also proposes a framework to quantify the varying degrees of robustness of Vision-Language Models (used as few-shot image classifiers) against spurious bias.
@@
Ranking Images by Spuriosity: One approach to detecting spurious bias in image classifiers is to rank images within their classes based on spuriosity, the degree to which common spurious cues are present. Prior work, such as HardImageNet [18] and Moayeri et al. [19], utilized deep neural features from an interpretable network combined with human supervision to identify these cues. In contrast, SpurLens eliminates the need for human supervision by leveraging object detectors to automatically rank images based on spurious cues. For example, in the case of a fire hydrant, our method can detect and rank images based on the presence of associated spurious cues, such as storm drains or sidewalks, without requiring manual annotations. Compared to prior work, SpurLens offers three key advantages: (1) a greater quantity of detected spurious cues, (2) improved interpretability via natural language descriptors, and (3) full automation without human involvement.
@@
Failures of Multimodal Systems: Some studies have introduced frameworks to automatically identify critical shortcomings of MLLMs [28, 29]. Tong et al. [29] highlight MLLMs’ struggles with basic visual understanding, attributing these issues to weaknesses in CLIP-based vision encoders. Conversely, Tong et al. [28] focuses on the language modality.
@@
Object hallucination: Object hallucination is an actively studied topic in MLLMs [2, 7, 14, 8, 38]. POPE [8] presents the first systematic study on object hallucination in MLLMs, providing an evaluation framework to quantify hallucination. Similarly, Leng et al. [7] provides additional evidence of object hallucination in MLLMs. Lovenia et al. [14] argues that hallucination should be distinguished from general incorrectness, introducing a separate benchmark for evaluating object hallucination in MLLMs. Hu et al. [2] proposes CIEM, an automatic pipeline that leverages an annotated image-text dataset along with an LLM to generate factual and contrastive question-answer pairs for evaluating hallucination in MLLMs. The authors also use CIEM to instruction-tune MLLMs in an attempt to mitigate hallucination. Additionally, Zhou et al. [38] introduces LURE, a post-hoc method that reduces object hallucination in MLLMs by reconstructing less hallucinatory descriptions. We emphasize that, although our evaluation framework follows POPE [8] and we adopt the evaluation metrics from Leng et al. [7], our study is not focused on evaluating object hallucination in MLLMs. The key difference between our work and the studies in this section is that we evaluate the reliance of MLLMs on spurious biases, rather than object hallucination itself.
@@ 3 Problem Setting
Figure 3: An overview of SpurLens. Left: GPT-4 is used to propose (and filter) spurious features $f_{i}$. OWLv2 annotates an image dataset with these spurious features, and its scores are used to rank the image dataset for each features. Right: for a given ranking, the top-$K$ and bottom-$K$ images are passed to an MLLM with a prompt. The responses are aggregated and used to compute the Spurious Gap.
We briefly present our theoretical framework for studying spurious bias in a multimodal setting. We adopt the common setup described in Ye et al. [35], where $\mathcal{X}$ and $\mathcal{Y}$ represent the vision and language modalities, respectively. Given an image input $x\in\mathcal{X}$ and a text input (prompt) $y\in\mathcal{Y}$, a MLLM learns the mapping $\mathcal{\phi:X\times Y\to O}$ such that $o=\mathcal{\phi}(x,y)$, where $o\in\mathcal{O}\subset\mathcal{Y}$ denotes the response.
To study spurious biases, for an object, we define $\mathcal{C\subset X}$ as the set of images that include target object $t$, and $\mathcal{S\subset X}$ as the set of images containing spurious cue $f$ associated with that object. We can partition $\mathcal{X}$ into 4 subsets: Object with Spurious Cue ($\mathcal{C\cap S}$), Object without Spurious Cue ($\mathcal{C\cap S^{\text{c}}}$), Spurious Cue without Object ($\mathcal{C^{\text{c}}\cap S}$), and Baseline (Object and Spurious Cue Free; $\mathcal{C^{\text{c}}\cap S^{\text{c}}}$).
[…]
Note that spurious attributes may also exist in the language modality. In our experiments, we minimized their impact by averaging over multiple simple binary prompts. Moreover, due to the flexibility of language and the extensive amount of text data, spurious attributes in the text modality often exhibit only weak correlations with specific responses, such as “yes” or “no” [35].
Similar to Leng et al. [7], we used two core metrics for our experiments: Perception Accuracy (PA) and Hallucination Rate (HR). Formally:
$\textbf{PA}:=p(o=\text{`yes'}\mid x\in\mathcal{C},y)$ (1)
$\textbf{HR}:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}},y)$ (2)
PA measures the model’s ability to accurately perceive objects that are present, while HR quantifies the hallucination of non-existent objects. Higher PA scores indicate better perception, whereas lower HR scores reflect greater robustness against object hallucinations. To study spurious bias, we analyze how PA and HR change with and without spurious cues for a given target object $t$. Formally, for object recognition we define the Spuriosity Gap as:
$\textbf{PA}_{s}$ $:=p(o=\text{`yes'}\mid x\in\mathcal{C\cap S},y)$ (3)
$\textbf{PA}_{c}$ $:=p(o=\text{`yes'}\mid x\in\mathcal{C\cap S^{\text{c}}},y)$ (4)
PA Gap $:=\textbf{PA}_{s}-\textbf{PA}_{c}$ (5)
Similarly, for object hallucination, we define the Spuriosity Gap as:
$\textbf{HR}_{s}$ $:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}\cap S},y)$ (6)
$\textbf{HR}_{c}$ $:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}\cap S^{\text{c}}},y)$ (7)
HR Gap $:=\textbf{HR}_{s}-\textbf{HR}_{c}$ (8)
A high Spuriosity Gap indicates that the MLLM is highly dependent on spurious cues to identify the presence of object $t$. Appendix B provides further theoretical intuition for these choices of metrics. We believe that modern MLLMs are susceptible to spurious correlations, and thus exhibit high PA and HR Gaps in many cases. We estimate these values by applying select MLLMs to images from all four partitions of $\mathcal{X}$, once they have been identified.
4 SpurLens: Automating Discovery of Spurious Cues
The spuriosity rankings in HardImageNet [18, 19] are constructed using spurious neural features extracted from the Salient ImageNet [27]. This process relies on human supervision to identify which features are spurious for each class. However, it is limited in scope, detecting only a few spurious features per class. Notably, for nearly two-thirds of ImageNet classes, no spurious cues were detected at all [19]. To enable the study of spurious correlations in MLLMs across a broader range of objects and datasets, we develop SpurLens, a pipeline designed to produce interpretable spuriosity rankings of images, from which the Spuriosity Gap can be estimated (outlined in Figure 3).
Suppose that, for a given MLLM $\mathcal{M}$ and target object $t$, we aim to identify which objects or image features act as spurious cues for $t$. Further, assume that we have access to a large dataset of images, denoted as $\{\mathcal{I}_{j}\}_{j=1}^{N}$, and that the presence of $t$ in each image is already annotated. If we are testing object recognition, all images must contain the target object $t$; if we are testing object hallucination, all images must not contain $t$.
Since we assume that the dataset provides ground-truth labels for the presence/absence of $t$, we do not use object detectors to determine whether the target object is present. This distinction is crucial because we believe that object detectors themselves are also susceptible to spurious correlations (using image encoders similar to the MLLMs we evaluate). Instead, we use object detectors solely for identifying spurious cues, ensuring that our analysis focuses on the MLLM’s behavior rather than the reliability of object detection models.
@@ 4 SpurLens: Automating Discovery of Spurious Cues
Proposing Spurious Features We use GPT-4 to generate list a list of objects or background elements that commonly appear in images of $t$. The number of features produced and their relation to $t$ can be easily adjusted to adapt to characteristics of the available dataset. We lemmatize each suggested object, remove duplicates, and remove features that share words with target object name. We then ask GPT-4 several filtering questions to ensure that the proposed objects match the qualifications for being spurious features. See Appendix A for details of these prompts.
@@
Works such as Leng et al. [7], Zhou et al. [38] identify spurious correlations through the frequent co-occurrence of objects in MLLM-generated image captions. Our method avoids this computational cost, and the easily-modifiable prompt structure may suggest a more diverse pool of potential spurious objects.
@@
Identifying Spurious Objects To identify the presence of these spurious features $f_{i}$ in the images $\mathcal{I}_{j}$, we use the OWLv2 open-set object detector [17]. For each image, we query OWLv2 with all potential spurious features and obtain several triplets consisting of a bounding box $b\in[0,1]^{4}$, label $f_{i}$, and confidence score $c\in[0,1]$. Let $\mathcal{O}(\mathcal{I}_{i})$ denote the set of such triplets produced by OWLv2 for image $\mathcal{I}_{i}$. We define the $f_{i}$-score of $\mathcal{I}_{j}$ as
@@
For each potential spurious feature $f_{i}$, we sort the images by $f_{i}$-score to obtain a ranking. Further discussion and a validation object detection results is provided in Appendix E.
Spuriosity Gaps For each ranking corresponding to feature $f_{i}$, let $\mathcal{U}_{t,f_{i}}^{+},\mathcal{U}_{t,f_{i}}^{-}\subset\{\mathcal{I}_{j}\}_{j=1}^{N}$ be the images with the $K$-highest and $K$-lowest $f_{i}$-scores respectively. For each of these images, we apply the model $\mathcal{M}$ paired with three prompts $p_{k}(t)$, $1\leq k\leq 3$ (see Appendix A for details). Each prompt asks $\mathcal{M}$ if it sees the target object $t$ in the image, and elicits a Yes/No response; we use three prompts to mitigate the bias due to word choice. We define PA (or similarly HR) of $\mathcal{M}$ on image $\mathcal{I}$ as:
[…]

### v2

# SpurLens: Automatic Detection of Spurious Cues in Multimodal LLMs

@@ Abstract
Unimodal vision models are known to rely on spurious correlations, but it remains unclear to what extent Multimodal Large Language Models (MLLMs) exhibit similar biases despite language supervision. In this paper, we investigate spurious bias in MLLMs and introduce SpurLens, a pipeline that leverages GPT-4 and open-set object detectors to automatically identify spurious visual cues without human supervision. Our findings reveal that spurious correlations cause two major failure modes in MLLMs: (1) over-reliance on spurious cues for object recognition, where removing these cues reduces accuracy, and (2) object hallucination, where spurious cues amplify the hallucination by over 10x. We investigate various MLLMs and datasets, and validate our findings with multiple robustness checks. Beyond diagnosing these failures, we explore potential mitigation strategies, such as prompt ensembling and reasoning-based prompting, and conduct ablation studies to examine the root causes of spurious bias in MLLMs. By exposing the persistence of spurious correlations, our study calls for more rigorous evaluation methods and mitigation strategies to enhance the reliability of MLLMs.
@@ 1 Introduction
Multimodal large language models (MLLMs) (Qwen2VL; llavanext; llama3; openai2024gpt4omini) have seen rapid advances in recent years. These models leverage the powerful capabilities of large language models (LLMs) (openai2024gpt4technicalreport; llama2) to process diverse modalities, such as images and text. They have demonstrated significant proficiency in tasks such as image perception, visual question answering, and instruction following.
Despite these advancements, MLLMs still exhibit critical visual shortcomings (multimon; eyeswideshut). One such failure is object hallucination (POPE; Ciem; NOPE; cursemultimodal), where MLLMs generate semantically coherent but factually incorrect content, falsely detecting objects that are not present in the input images, as exemplified in Figure 1. We hypothesize that many of these failures stem from a well-known robustness issue in deep learning models: spurious bias – the tendency to rely on non-essential input attributes rather than truly recognizing the target object (MMSpuBench). While spurious correlation reliance has been well-documented in single-modality image classifiers, the extent to which it persists in MLLMs remains unclear.
Understanding spurious correlations is crucial because they can lead to systematic failures. Consider the case of a fire hydrant (Figure LABEL:fig:intro_eval). When a fire hydrant appears in a street scene, Llama-3.2 (llama3) correctly recognizes it 96% of the time. However, when placed in unusual contexts, such as a warehouse, accuracy drops to 83%, suggesting an over-reliance on contextual cues rather than the object itself. Conversely, when shown a street scene without a fire hydrant, the model hallucinates its presence 7 times more often than when given a random image without a hydrant, such as an indoor environment lacking street-like features.
Our results reveal that MLLMs over-rely on spurious cues, leading to two major failures. The first is object recognition failures: when spurious cues are removed, recognition performance significantly drops. For example, GPT-4o-mini (openai2024gpt4omini) sees a 20.4% accuracy drop on COCO (coco), from 88.0% to 67.6% (Table 1). The second is object hallucination: spurious cues trigger hallucinations, causing models to misidentify objects. For GPT-4o-mini, hallucination rates increase 26-fold when spurious cues are present (Table 2). We illustrate some of these failures in Figure 1.
Object: Tie
Missing Spurious Cue: Suit
@@
Object: Toaster
Missing Spurious Cue: Kitchen Counter
@@
There is a black coffee maker on the countertop, but I don’t see a toaster in the image. The area appears to have kitchen cabinets and …
Object: Balance Beam
Spurious Cue: Rope
@@
Object: Ruffed Grouse
Spurious Cue: Moss
@@
Figure 1: Some failures of GPT-4o-mini identified by SpurLens. (Left) The model fails to recognize objects in the absence of spurious cues. (Right) Spurious cues trigger hallucinations.
@@
SpurLens offers three key advantages over prior work. First, SpurLens enables the detection of a broader range of spurious cues as it is not limited to predefined features. Second, it provides natural language descriptors for detected cues, offering insight into model biases. Finally, our method requires no human supervision, making it scalable and adaptable across datasets and objects. We validate the reliability and robustness of SpurLens through human studies and analysis of its components. Its outputs align closely with human judgments and remain consistent when using different language models to propose spurious cues, highlighting its generality and reproducibility.
@@
While SpurLens allows us to systematically analyze spurious correlations in MLLMs, some key questions remain: from where do these biases originate, and can they be mitigated? To address these, we conduct a series of ablation studies designed to isolate different factors contributing to spurious bias. First, we investigate whether spurious bias persists even when all visual evidence of an object is artificially removed by dropping the corresponding visual tokens. We find that even when target objects have been excised from images, MLLMs tend to hallucinate their presence at a high rate. Next, we explore whether alternative prompting strategies can mitigate spurious reliance. We find that prompt ensembling and reasoning-based prompting offer slight situational improvements but do not significantly reduce spurious bias. Even informing the MLLM of spurious features in the prompt does not meaningfully improve performance, indicating that the spurious biases identified by SpurLens are fundamental issues. Finally, we examine whether spurious correlations can be measured purely from the MLLM image embeddings. We find that the vision encoder alone also exhibits spurious biases, reinforcing that the issue extends beyond multimodal fusion errors. These findings indicate that spurious bias is a deeply rooted problem in MLLMs and that simple mitigation strategies are insufficient to fully address it.
@@ 2 Related Work
Spurious Correlation: Spurious correlations have been extensively studied in the context of deep neural network classifiers (e.g., ViT (ViT)), with various approaches proposed to detect and mitigate the issue (groupdro; lastlayer; dac). However, these studies primarily focus on uni-modal settings (image classification tasks). Some research (soberclip; RaVL; kim2023exposing) has explored spurious correlations in CLIP (clip), framing the problem in terms of zero-shot performance across vision and language modalities. MMSpuBench introduces a visual question answering (VQA) benchmark designed to evaluate MLLMs’ reliance on spurious correlations using open-source image datasets. FewSTAB also proposes a framework to quantify the varying degrees of robustness of Vision-Language Models (used as few-shot image classifiers) against spurious bias.
Interpretable Spurious Biases: Some works investigate language-interpretable and semi-automatic detection of spurious biases in vision models. eyuboglu2022dominodiscoveringsystematicerrors and zhang2023diagnosingrectifyingvisionmodels both identify interpretable high-error slices using cross-modal embeddings; however, the former requires additional feature annotations on each sample, while the latter requires human selection of potential text attributes. wiles2023discoveringbugsvisionmodels uses clustering of failure cases and image captioning to identify interpretable spurious features; however, they use image generation models to produce synthetic datasets for iterative refinement. In contrast to these methods, SpurLens does not require human supervision, and consistently identifies strong failures in real-world image datasets.
@@
Ranking Images by Spuriosity: One approach to detecting spurious bias in image classifiers is to rank images within their classes based on spuriosity, the degree to which common spurious cues are present. Prior work, such as HardImageNet (hardimagenet) and moayeri2023spuriosity, utilized deep neural features from an interpretable network combined with human supervision to identify these cues. In contrast, SpurLens eliminates the need for human supervision by leveraging object detectors to automatically rank images based on spurious cues.
@@
Failures of Multimodal Systems: Some studies have introduced frameworks to automatically identify critical shortcomings of MLLMs (multimon; eyeswideshut). eyeswideshut highlight MLLMs’ struggles with basic visual understanding, attributing these issues to weaknesses in CLIP-based vision encoders. Conversely, multimon focuses on the language modality.
@@
Object Hallucination: Object hallucination is an actively studied topic in MLLMs (Ciem; cursemultimodal; NOPE; POPE; zhou2023analyzing). POPE (POPE) presents the first systematic study on object hallucination in MLLMs, providing an evaluation framework to quantify hallucination. Similarly, cursemultimodal provides additional evidence of object hallucination in MLLMs. NOPE argues that hallucination should be distinguished from general incorrectness, introducing a separate benchmark for evaluating object hallucination in MLLMs. Ciem proposes CIEM, an automatic pipeline that leverages an annotated image-text dataset along with an LLM to generate factual and contrastive question-answer pairs for evaluating hallucination in MLLMs. The authors also use CIEM to instruction-tune MLLMs in an attempt to mitigate hallucination. Additionally, zhou2023analyzing introduces LURE, a post-hoc method that reduces object hallucination in MLLMs by reconstructing less hallucinatory descriptions. We emphasize that, although our evaluation framework follows POPE (POPE) and we adopt the evaluation metrics from cursemultimodal, our study is not focused on evaluating object hallucination in MLLMs. The key difference between our work and the studies in this section is that we evaluate the reliance of MLLMs on spurious biases, rather than object hallucination itself.
@@ 3 Problem Setting
We briefly present our theoretical framework for studying spurious bias in a multimodal setting. We adopt the common setup described in MMSpuBench, where $\mathcal{X}$ and $\mathcal{Y}$ represent the vision and language modalities, respectively. Given an image input $x\in\mathcal{X}$ and a text prompt $y\in\mathcal{Y}$, a MLLM learns the mapping $\mathcal{\phi:X\times Y\to O}$ such that $o=\mathcal{\phi}(x,y)$, where $o\in\mathcal{O}\subset\mathcal{Y}$ denotes the response.
To study spurious biases, for an object, we define $\mathcal{C\subset X}$ as the set of images that include target object $t$, and $\mathcal{S\subset X}$ as the set of images containing spurious cue $f$ associated with that object. We can partition $\mathcal{X}$ into 4 subsets: Object with Spurious Cue ($\mathcal{C\cap S}$), Object without Spurious Cue ($\mathcal{C\cap S^{\text{c}}}$), Spurious Cue without Object ($\mathcal{C^{\text{c}}\cap S}$), and Baseline (Object and Spurious Cue Free; $\mathcal{C^{\text{c}}\cap S^{\text{c}}}$). For example, if t=t=“fire hydrant” and f=f=“road”, $\mathcal{C}$ is all images including a fire hydrant, and $\mathcal{S}$ is all images including a road. For prompt y=y=“Is there a fire hydrant in the image?", the correct response $o$ is “yes” for $x\in\mathcal{C}$ and “no” for $x\not\in\mathcal{C}$. Examples of all four types of images are found in Figure LABEL:fig:intro_eval.
Note that spurious attributes may also exist in the language modality. In our experiments, we minimize their impact by averaging over multiple simple binary prompts. Moreover, due to the flexibility of language and the extensive amount of text data, spurious attributes in the text modality often exhibit only weak correlations with specific responses, such as “yes” or “no” (MMSpuBench).
Similar to cursemultimodal, we used two core metrics for our experiments: Perception Accuracy (PA) and Hallucination Rate (HR). Formally:
$\textbf{PA}:=p(o=\text{`yes'}\mid x\in\mathcal{C},y),\;\;\;\textbf{HR}:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}},y)$ (1)
PA measures the model’s ability to accurately perceive objects that are present, and HR quantifies the hallucination of non-existent objects. Higher PA scores indicate better perception, while lower HR scores reflect greater robustness against object hallucinations. To study spurious bias, we analyze how PA and HR change with and without spurious cues for a given target object $t$. Formally, for object recognition the Spurious Gap (moayeri2023spuriosity) is defined as $\textbf{PA Gap}:=\textbf{PA}_{s}-\textbf{PA}_{c}$, where:
$\textbf{PA}_{s}:=p(o=\text{`yes'}\mid x\in\mathcal{C\cap S},y)$ $\textbf{PA}_{c}:=p(o=\text{`yes'}\mid x\in\mathcal{C\cap S^{\text{c}}},y)$ (2)
Similarly, for object hallucination, the Spurious Gap is defined as $\textbf{HR Gap}:=\textbf{HR}_{s}-\textbf{HR}_{c}$, where:
$\textbf{HR}_{s}:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}\cap S},y)$ $\textbf{HR}_{c}:=p(o=\text{`yes'}\mid x\in\mathcal{C^{\text{c}}\cap S^{\text{c}}},y)$ (3)
A high Spurious Gap indicates that the MLLM is highly dependent on spurious cues to identify the presence of object $t$. Appendix B provides further theoretical intuition for these choices of metrics. We believe that modern MLLMs are susceptible to spurious correlations, and thus exhibit high PA and HR Gaps in many cases. We estimate these values by applying selected MLLMs to images from all four partitions of $\mathcal{X}$, once they have been identified.
4 SpurLens: Automating Discovery of Spurious Cues
The spuriosity rankings in HardImageNet (hardimagenet; moayeri2023spuriosity) are constructed using neural features extracted from Salient ImageNet (salientimagenet). This process relies on human supervision to identify spurious features for each class. However, it is limited in scope, detecting only a few spurious features per class. Notably, for nearly two-thirds of ImageNet classes, no spurious cues were detected at all (moayeri2023spuriosity). To study spurious correlations in MLLMs across a broader range of objects and datasets, we develop SpurLens, a pipeline designed to produce interpretable spuriosity rankings of images, from which the Spurious Gap can be estimated (outlined in Figure 2).
Suppose that, for a given MLLM $\mathcal{M}$ and target object $t$, we aim to identify spurious cues for $t$. Assume that we have access to a large dataset of images, denoted as $\{\mathcal{I}_{j}\}_{j=1}^{N}$, and that the presence of $t$ in each image is known; if we are testing object recognition, all images must contain the target object $t$; if we are testing object hallucination, all images must not contain $t$.
@@ 4 SpurLens: Automating Discovery of Spurious Cues
Figure 2: An overview of SpurLens. Left: GPT-4 is used to propose (and filter) spurious features $f_{i}$. OWLv2 annotates an image dataset with these spurious features, and its scores are used to rank the image dataset for each feature. Right: for a given ranking, the top-$K$ and bottom-$K$ images are passed to an MLLM with a prompt. The responses are aggregated to compute the Spurious Gap.
@@
Proposing Spurious Features We use GPT-4 to generate a list of objects or background elements that commonly appear in images of $t$. We lemmatize each suggested object, remove duplicates, and remove features that share words with target object name. We then ask GPT-4 several filtering questions to ensure that the proposed objects match the qualifications for being spurious features, and to preemptively eliminate objects that may be challenging for the object detector (e.g., features that are not easily detectable such as “sunlight”, or features that are part of the corresponding target object such as “screen” for “laptop”). Note that this filtering reduces the set of potential spurious features in exchange for improved robustness in object detection. See Appendix A for details of these prompts.
@@
Works such as cursemultimodal; zhou2023analyzing; kim2024discoveringmitigatingvisualbiases identify spurious correlations through the frequent co-occurrence of objects in MLLM-generated image captions. Our method avoids this computational cost, and the easily-modifiable prompt structure may suggest a more diverse pool of potential spurious objects.
@@
Identifying Spurious Objects To identify the presence of these spurious features $f_{i}$ in the images $\mathcal{I}_{j}$, we use the OWLv2 open-set object detector owlv2. For each image, we query OWLv2 with all potential spurious features and obtain several triplets consisting of a bounding box $b\in[0,1]^{4}$, label $f_{i}$, and confidence score $c\in[0,1]$. Let $\mathcal{O}(\mathcal{I}_{i})$ denote the set of such triplets produced by OWLv2 for image $\mathcal{I}_{i}$. We define the $f_{i}$-score of $\mathcal{I}_{j}$ as
@@
For each potential spurious feature $f_{i}$, we sort the images by $f_{i}$-score to obtain a ranking. Validation of object detection results is provided in Appendix F.
Spurious Gaps For each ranking corresponding to feature $f_{i}$, let $\mathcal{U}_{t,f_{i}}^{+},\mathcal{U}_{t,f_{i}}^{-}\subset\{\mathcal{I}_{j}\}_{j=1}^{N}$ be the images with the $K$-highest and $K$-lowest $f_{i}$-scores respectively. For each of these images, we apply the model $\mathcal{M}$ paired with three prompts $p_{k}(t)$, $1\leq k\leq 3$ (see Appendix A for details). Each prompt asks $\mathcal{M}$ if it sees the target object $t$ in the image, and elicits a Yes/No response; we use three prompts to mitigate the bias due to word choice. We define PA (or similarly HR) of $\mathcal{M}$ on image $\mathcal{I}$ as:
$\textbf{PA}(\mathcal{M},\mathcal{I},t)=\frac{1}{3}\sum\nolimits_{k=1}^{3}\mathds{1}\left(\mathcal{M}(\mathcal{I},p_{k}(t))=\text{``Yes"}\right)$ (5)